# The Dark Side of NFTs: A Large-Scale Empirical Study of Wash Trading

### v1

@@ IV-C Dataset overview
Table III demonstrates the final look of our dataset, and we make some preliminary calculations on it, e.g., holders of each NFT change on average 3.2 times each year, indicating that the NFT market does not enjoy high liquidity. Therefore, consecutive TXNs in a short period are abnormal and worth being concerned about. Also, the selected block TXNs and ERC-20 token TXNs contain ETH and ERC-20 token transfer information between each NFT sale event’s from and to, exposing that the related addresses generate more behaviors before/after the NFT TXNs, as there is no direct transfer of ETH or ERC-20 token between two trading addresses for a common deal on OpenSea, except for the auction, using WETH as a payment token.
@@
TABLE III: The final look of our dataset applied to the identification algorithm.
V WASH TRADING/TRADERS IDENTIFICATION
Figure 2: The workflow of wash trading and wash trader identification
@@ V WASH TRADING/TRADERS IDENTIFICATION
V-A Approach on Round-trip Trading
V-A1 Approach design
@@
  • Whether a transfer event contains trading information is uncertain. As we mentioned in IV-B3, it is possible that a transfer event is essentially a real NFT trading. To extend the detection, we have included transfer events in our algorithm.
V-A2 Time window
@@
V-A3 Cycles finding
[…]
V-A4 Wash trading confirmation
@@
V-B Approach on Unprofitable Trading
@@
V-C Approach on Hidden Trading
@@
V-D Detection of wash trader
The probability of addresses appearing simultaneously as a pair or a group provides the reference for wash trader detection. We utilize FP-Growth  [19], an association-rule mining algorithm, on addresses related to Round-trip Trading, Unprofitable Trading, and Hidden Trading, to find the suspicious address pairs. We define the tuple R: (from, to) for Round-trip Trading, U: (from, to) for Unprofitable Trading, to represent the wash trading pairs. Considering the colluding relationship is not displayed by each sale event of Hidden Trading, we include every address of the continuous private trading, i.e., the tuple H: (address1, address2, address3, …). Given a set of addresses, frequent pattern mining can find all the itemsets with a frequency greater than the support. The matrix $[R_{all};U_{all};H_{all}]^{T}$ is the final input for FP-Growth, where the elements for each row vector come from the tuples of R, U, or H. If the address pair/group’s occurrence frequency exceeds the support, the address pair/group is detected as a wash trading pair/group. Also, we combine all wash trading pairs with at least one common address/TXN between each other as a wash trading group.
VI EXPERIMENTAL RESULTS
@@
TABLE IV: Top five collections by the number of wash trading behaviors
VI-A Results
VI-A1 Results overview
We identify 824 transfer events and 5,330 sale events related to NFT wash trading, accounting for \$8,857,070.41, which is around 0.12% of the total trading amount (\$7,216,023,387.61). It suggests that the financial impact of NFT wash trading on popular collections is relatively small and overrated. However, as shown in Table IV that displays the top five collections with their market value ranked by the number of wash trading behaviors, OG:Crystal with lagging ranking is wash traded most, far exceeding the other collections. It indicates the harm of wash trading inflicted on the NFT market might not be fully exposed through trading volume, inspiring us to explore broader aspects of its impact.
VI-A2 Results of Round-trip Trading
@@
VI-A3 Results of Unprofitable Trading
@@
VI-A4 Results of Hidden Trading
@@
Figure 3: Obvious evidence in visualization for wash trading groups.
VI-A5 Results of wash trading pairs/groups
@@
Figure 4: The trend for the number of events related to wash trading, excluding OG:Crystal.
VI-A6 Trend of wash trading events.
We investigate the trend for sale/transfer events related to NFT wash trading. Particularly, we exclude OG:Crystal to avoid its extremely leading role in our trend assessment. Figure 4 shows the change in the number of wash trading events of our results from 2021-06-28T15:00:34Z to 2022-06-20:18:28:10Z. The events have grown and become stable since June 2021 and reach their peak on January 11, 2022. It indicates that wash trading has been used consciously after June 2021. For the peak, it is mostly because of the launch of LooksRare [37], an NFT marketplace with an incentive reward plan that encourages frequent TXNs [38] and thus spawns a bunch of wash trading behaviors [15, 39, 40] (See VII-A for details).
Figure 5: Time interval differences when transitioning to the next non-wash-traded time window between wash-traded and non-wash-traded time windows
VI-A7 Market liquidity.
@@
VII FINDINGS FROM MULTIPLE DIMENSIONS
@@ VII-B Profitability
We investigate the possibility for wash traders to profit from benign users without considering incentive rewards from NFT marketplaces through the results of Round-trip Trading. In other words, we assess if wash traders can offset the service fees (charged by NFT marketplaces for their service) by reselling the NFTs. Our analysis focuses on a scenario where a non-wash-traded time window follows a wash-traded time window. At the same time, the total service fees from round-trip sale events in the previous time window are less than or equal to the first sale event in the latter time window. To strengthen our analysis, we adopt OpenSea’s service fee rate (2.5%), the highest among major NFT marketplaces [41]. Our analysis identifies 1,752 wash-traded time windows that satisfy the scenario. Among them, wash traders in around 60% of cases achieve profitability when jumping to the next non-wash-traded time window. Figure 7 shows the distribution of all wash-traded time windows regarding gain or loss, where most are concentrated around 0, indicating that most wash traders do not suffer significant losses after deducting the service fees. For example, in the case of OG:Crystal #895, two traders (0x89a09c and 0x0B7742) traded the NFT six times with prices ranging from 0.06 ETH to 0.09 ETH during the first time window. The NFT was finally sold for 1 ETH in the first sale event of the second time window, resulting in a profit of 0.99 ETH for wash traders.
Figure 6: The distribution of all wash-traded time windows in terms of gain or loss
Figure 6: The distribution of all wash-traded time windows in terms of gain or loss
Figure 7: The difference between the last trading price in wash-traded time window with each trading price for the following 1st to 7th users
@@
VII-C NFT project design
Figure 8: The evolution of OG:Crystal #1861 from the 5th generation to its final form through Round-trip Trading [42]
As Table IV shown, OG:Crystal is the collection with the most wash trading behaviors totaling 9,412 times, accounting for \$1,031,442.16. However, the main purpose of wash trading on OG:Crystal is not to inflate trading volume but to enhance the appearance of the NFTs. The design and working mechanism of OG:Crystal is responsible for this [43, 27]: each OG:Crystal transforms and evolves with each purchase made by a new collector. The appearance of each piece is generated based on the properties of the owner’s crypto wallet. Specifically, each OG:Crystal will be ‘locked’, meaning that new TXNs will not affect its shape, structure, or rarity, two months after the initial sale date or after it has grown by seven generations through multiple trading. For example, Figure 8 shows the evolution of OG:Crystal #1861 from 5th generation to 7th generation through Round-trip Trading between 0x9FFFB8 and 0xB05Cb1. The overview result on OG:Crystal is more convincing for verifying its working mechanism. 1) Wash traders of OG:Crystal generally appear as a pair, and the average number of walks formed by TXNs of each NFT is nearly 35, meaning that they may resell the NFT to each other around six times. 2) Wash trading does not exist after the NFT is ‘locked’ in 95.73% cases. Both findings suggest that the design of OG:Crystal can lead to wash trading problem, turning the project’s vision that creates a public work of art for all users into personal wash trading activities.
On the contrary, Kaiju Kingz is a collection that exhibits minimal wash trading behavior, with a deliberate feature of its design [44]. Owners of Kaiju Kingz’s NFTs can generate 5 RWASTE per day, the circulating token for Kaiju Kingz used for breeding. They can create a so-called Kaiju Kingz baby (a new NFT) with 750 RWASTE. In this sense, owners will be more inclined to hold the NFTs for a longer period of time.
@@ VII-C NFT project design
VII-D Payment token
We haven’t discovered various ERC-20 tokens involved in Unprofitable Trading based on the results, and the main circulating ERC-20 token involved is WETH, accounting for 99.3%. WETH enables users to submit pre-authorized bids that will be automatically fulfilled later without the bidder’s permission. In our results, WETH mainly appears in two places. 1) NFT holders invoke the smart contract, Wrapped Ether, to transfer WETH to NFT bidders, and afterward, NFT holders accept NFT bidders’ offer to obtain WETH back through the auction. For example, Table V demonstrates the period of how 0xcF5e38 and 0xe67753 execute Unprofitable Trading: before the bidder 0xe67753 made an offer of 0.1 WETH to buy Chibi Dino #5723, 0xcF5e38 transferred total 0.1 + 0.07 + 0.3 + 0.47 = 0.94 WETH to 0xe67753. Two minutes after, 0xcF5e38 accepted the offer. It is a complete self-directed purchase that contributes to false trading volume. 2) Another occurrence shows in the situation that there is another NFT sale event with the opposite from and to as the current one 80 minutes before. The WETH was transferred from from to to through the previous bidding auction and is transferred back through the current one. It is Round-trip Trading with multiple bids won, e.g., 0xeB1543 and 0xB893AE accept each other’s offer respectively on OG:Crystal #5609 and OG:Crystal #6255 within one minute, and their offer prices are the same, i.e., 0.13 WETH.
@@ VII-D Payment token
TABLE V: An example of Unprofitable Trading using ERC-20 token: Chibi Dino #5723
[…]
@@
VII-E User behavior
Figure 9: The total number of ETH transfer events at different times before and after an NFT sale event.
In Unprofitable Trading, sellers collude with buyers to create unprofitable TXNs. For example, just 4 seconds before the sale event 0x1a4521, 0x4EDE98 completed a transfer (0.96 ETH) fully sufficient to support 0xE1f14d in buying Animeta #4302 (0.435 ETH) in a normal TXN 0x536fa8. To understand to what extent ETH transfer helps with each sale event, we define $M_{r}$ as the total amount received by sellers, $M_{t}$ as the total amount returned by sellers, and $M_{s}$ as the related NFT trading price. We then adopt Pearson correlation coefficient [45] to measure the strength of the linear relationship between ($M_{r}$, $M_{s}$) and ($M_{t}$, $M_{s}$). The results are 0.7999 for ($M_{r}$, $M_{s}$) and 0.9840 for ($M_{t}$, $M_{s}$), indicating that the amount of each ETH transfer is similar to the NFT trading price. Furthermore, we count the number of ETH transfer events detected under a finer granularity (i.e., 60 seconds). Figure 9 shows that the closer the time point before/after a sale event is, the more ETH transfer behaviors happen. Also, wash traders make the most frequent ETH transfer between two and four minutes before/after a sale event, after which the number gradually decreases. The respective high price and time correlations prove that these ETH transfers are directly for the NFT TXNs.
@@ VII-F NFT ecosystem
TABLE VI: Statistical results of decreased wash trading behavior
To better understand the impact of wash trading pairs/groups on NFT wash trading, we identify all 10,594 TXNs related to these address pairs and exclude their trading volume. For instance, if there is a flagged TXN in Hidden Trading, we do not calculate the trading volume of all consecutive private trading. As shown in Table VI, the results indicate a decrease of around 30% in wash trading amount for round-trip and Hidden Trading, while Unprofitable Trading experiences a significant decline, indicating the prevalence of wash traders in Unprofitable Trading. Interestingly, the addresses involved in wash trading pairs/groups accounted for only 0.082% of the total, meaning that a small group of users could cause considerable wash trading, and sizable wash trading behaviors could be eliminated by identifying and flagging them. It suggests that to avoid wash trading better and alert the users, NFT marketplaces do not need to keep all the trading addresses under governance but mark a small proportion of malicious accounts in event sequences.
@@
VIII DISCUSSION
VIII-A Threats to validity
Internal validity. 1) Adjusting the support for FP-Growth may affect the experimental results, e.g., the number of wash trading pairs/groups would increase if it is set loosely. Moreover, the support should be reconsidered for different datasets. To ensure the reliability of our results, we set a strict value, i.e., address pairs will be considered suspicious only when they participate in wash trading at least twelve times. 2) Our method may ignore unknown wash trading. We have not yet investigated more complex trading networks, e.g., incorporating internal TXNs to understand the intention of NFT traders from a more underlying perspective and thus reveal more NFT wash trading patterns. However, we are still the most comprehensive study on NFT wash trading.
External validity. The current tools (e.g., OpenSea API) for collecting NFT events have shortcomings, such as mislabeling transfer/sale/minted events. Nevertheless, we raise and address three issues related to our datasets and make public a guaranteed dataset for researchers.
VIII-B Implications for NFT Marketplace

### v2

@@ 4.4. Dataset overview
Table 2 demonstrates the final look of our dataset, and we make some preliminary calculations on it, e.g., holders of each NFT change on average 3.2 times each year, indicating that the NFT market does not enjoy high liquidity. Therefore, consecutive TXNs in a short period are abnormal and worth being concerned about. Also, the selected block TXNs and ERC-20 token TXNs contain ETH and ERC-20 token transfer information between each NFT sale event’s from and to, exposing that the related addresses generate more behaviors before/after the NFT TXNs, as there is no direct transfer of ETH or ERC-20 token between two trading addresses for a common deal on OpenSea, except for the auction, using WETH as a payment token.
@@
Table 2. The final look of our dataset applied to the identification algorithm.
5. DETECTION
Figure 2. The workflow of wash trading and wash trader identification
@@ 5. DETECTION
5.1. Approach on Round-trip Trading
5.1.1. Approach design
@@
  • Whether a transfer event contains trading information is uncertain. It is possible that a transfer event is essentially a real NFT trading. To extend the detection, we have included transfer events in our algorithm.
5.1.2. Time window
@@
(1) $\textit{ATI}=\frac{\textit{timestamp}_{\textit{end}}-\textit{timestamp}_{\textit{start}}}{\textit{num}}$
@@ 5.1.3. Cycles finding
5.1.4. Wash trading confirmation
@@
5.2. Approach on Unprofitable Trading
@@
5.3. Approach on Hidden Trading
@@
5.4. Detection of wash trader
The probability of addresses appearing simultaneously as a pair or a group provides the reference for wash trader detection. We utilize FP-Growth  (Han et al., 2004), an association-rule mining algorithm, on addresses related to Round-trip Trading, Unprofitable Trading, and Hidden Trading, to find the suspicious address pairs. We define the tuple R: (from, to) for Round-trip Trading, U: (from, to) for Unprofitable Trading, to represent the wash trading pairs. Considering the colluding relationship is not displayed by each sale event of Hidden Trading, we include every address of the continuous private trading, i.e., the tuple H: (address1, address2, address3, …). Given a set of addresses, frequent pattern mining can find all the itemsets with a frequency greater than the support. The matrix $[R_{all};U_{all};H_{all}]^{T}$ is the final input for FP-Growth, where the elements for each row vector come from the tuples of R, U, or H. If the address pair/group’s occurrence frequency exceeds the support, the address pair/group is detected as a wash trading pair/group. Also, we combine all wash trading pairs with at least one common address/TXN between each other as a wash trading group.
6. EXPERIMENTAL RESULTS
@@
Table 3. Top five collections by the number of wash trading behaviors
6.1. Results
6.1.1. Results overview
We identify 824 transfer events and 5,330 sale events related to NFT wash trading, accounting for \$8,857,070.41, which is around 0.12% of the total trading amount (\$7,216,023,387.61). It suggests that the financial impact of NFT wash trading on popular collections is relatively small and overrated. However, as shown in Table 3 that displays the top five collections with their market value ranked by the number of wash trading behaviors, OG:Crystal with lagging ranking is wash traded most, far exceeding the other collections. It indicates the harm of wash trading inflicted on the NFT market might not be fully exposed through trading volume, inspiring us to explore broader aspects of its impact.
6.1.2. Results of Round-trip Trading
@@
6.1.3. Results of Unprofitable Trading
@@
6.1.4. Results of Hidden Trading
@@
Figure 3. Obvious evidence in visualization for wash trading groups.
6.1.5. Results of wash trading pairs/groups
@@
Figure 4. The trend for the number of events related to wash trading, excluding OG:Crystal.
6.1.6. Trend of wash trading events.
We investigate the trend for sale/transfer events related to NFT wash trading. Particularly, we exclude OG:Crystal to avoid its extremely leading role in our trend assessment. Figure 4 shows the change in the number of wash trading events of our results from 2021-06-28T15:00:34Z to 2022-06-20:18:28:10Z. The events have grown and become stable since June 2021 and reach their peak on January 11, 2022. It indicates that wash trading has been used consciously after June 2021. For the peak, it is mostly because of the launch of LooksRare (off, 2023), an NFT marketplace with an incentive reward plan that encourages frequent TXNs (LooksRare, 2023) and thus spawns a bunch of wash trading behaviors (Serneels, 2023; La Morgia et al., 2022; Cho et al., 2023) (See 7.1 for details).
Figure 5. Time interval differences when transitioning to the next non-wash-traded time window between wash-traded and non-wash-traded time windows
6.1.7. Market liquidity.
@@
7. FINDINGS FROM MULTIPLE DIMENSIONS
@@ 7.2. Profitability
We investigate the possibility for wash traders to profit from benign users without considering incentive rewards from NFT marketplaces through the results of Round-trip Trading. In other words, we assess if wash traders can offset the service fees (charged by NFT marketplaces for their service) by reselling the NFTs. Our analysis focuses on a scenario where a non-wash-traded time window follows a wash-traded time window. At the same time, the total service fees from round-trip sale events in the previous time window are less than or equal to the first sale event in the latter time window. To strengthen our analysis, we adopt OpenSea’s service fee rate (2.5%), the highest among major NFT marketplaces. Our analysis identifies 1,752 wash-traded time windows that satisfy the scenario. Among them, wash traders in around 60% of cases achieve profitability when jumping to the next non-wash-traded time window. Figure 7 shows the distribution of all wash-traded time windows regarding gain or loss, where most are concentrated around 0, indicating that most wash traders do not suffer significant losses after deducting the service fees. For example, in the case of OG:Crystal #895, two traders (0x89a09c and 0x0B7742) traded the NFT six times with prices ranging from 0.06 ETH to 0.09 ETH during the first time window. The NFT was finally sold for 1 ETH in the first sale event of the second time window, resulting in a profit of 0.99 ETH for wash traders.
Figure 6. The distribution of all wash-traded time windows in terms of gain or loss
Figure 7. The difference between the last trading price in wash-traded time window with each trading price for the following 1st to 7th users
@@
7.3. NFT project design
Figure 8. The evolution of OG:Crystal #1861 from the 5th generation to its final form through Round-trip Trading (Reef, 2022)
As Table 3 shown, OG:Crystal is the collection with the most wash trading behaviors totaling 9,412 times, accounting for \$1,031,442.16. However, the main purpose of wash trading on OG:Crystal is not to inflate trading volume but to enhance the appearance of the NFTs. The design and working mechanism of OG:Crystal is responsible for this (ogw, 2023; La Morgia et al., [n. d.]): each OG:Crystal transforms and evolves with each purchase made by a new collector. The appearance of each piece is generated based on the properties of the owner’s crypto wallet. Specifically, each OG:Crystal will be ‘locked’, meaning that new TXNs will not affect its shape, structure, or rarity, two months after the initial sale date or after it has grown by seven generations through multiple trading. For example, Figure 8 shows the evolution of OG:Crystal #1861 from 5th generation to 7th generation through Round-trip Trading between 0x9FFFB8 and 0xB05Cb1. The overview result on OG:Crystal is more convincing for verifying its working mechanism. 1) Wash traders of OG:Crystal generally appear as a pair, and the average number of walks formed by TXNs of each NFT is nearly 35, meaning that they may resell the NFT to each other around six times. 2) Wash trading does not exist after the NFT is ‘locked’ in 95.73% cases. Both findings suggest that the design of OG:Crystal can lead to wash trading problem, turning the project’s vision that creates a public work of art for all users into personal wash trading activities.
On the contrary, Kaiju Kingz is a collection that exhibits minimal wash trading behavior, with a deliberate feature of its design (KaijuKingz, 2023). Owners of Kaiju Kingz’s NFTs can generate 5 RWASTE per day, the circulating token for Kaiju Kingz used for breeding. They can create a so-called Kaiju Kingz baby (a new NFT) with 750 RWASTE. In this sense, owners will be more inclined to hold the NFTs for a longer period of time.
@@ 7.3. NFT project design
7.4. Payment token
We haven’t discovered various ERC-20 tokens involved in Unprofitable Trading based on the results, and the main circulating ERC-20 token involved is WETH, accounting for 99.3%. WETH enables users to submit pre-authorized bids that will be automatically fulfilled later without the bidder’s permission. In our results, WETH mainly appears in two places. 1) NFT holders invoke the smart contract, Wrapped Ether, to transfer WETH to NFT bidders, and afterward, NFT holders accept NFT bidders’ offer to obtain WETH back through the auction. For example, Table 4 demonstrates the period of how 0xcF5e38 and 0xe67753 execute Unprofitable Trading: before the bidder 0xe67753 made an offer of 0.1 WETH to buy Chibi Dino #5723, 0xcF5e38 transferred total 0.1 + 0.07 + 0.3 + 0.47 = 0.94 WETH to 0xe67753. Two minutes after, 0xcF5e38 accepted the offer. It is a complete self-directed purchase that contributes to false trading volume. 2) Another occurrence shows in the situation that there is another NFT sale event with the opposite from and to as the current one 80 minutes before. The WETH was transferred from from to to through the previous bidding auction and is transferred back through the current one. It is Round-trip Trading with multiple bids won, e.g., 0xeB1543 and 0xB893AE accept each other’s offer respectively on OG:Crystal #5609 and OG:Crystal #6255 within one minute, and their offer prices are the same, i.e., 0.13 WETH.
@@ 7.4. Payment token
Table 4. An example of Unprofitable Trading using ERC-20 token: Chibi Dino #5723
During the auction process, we notice more obvious evidence of address collusion regarding WETH, e.g., some auction is essential almost free transfer of NFT, making the NFT’s historical price curve sag seriously and creating a huge deviation from the market reference value. For example, the accepted offer price for dotdotdot #2663 on December 30, 2021, is less than 0.0001 WETH, while its previous trading price is 3 ETH. The NFT holder and bidder know each other for such abnormal behavior, and their user names all start with blitmonk. To explore the observation, we define $\textit{PF}=\frac{\textit{Price}{\textit{current}}}{\textit{Price}{\textit{latest}}}$ to represent the trading price fluctuations, where we start counting from 0 for each NFT’s all sale events, i.e., the 0thPF is the price fluctuation between an NFT’s first and second sale events. To demonstrate server price sag, we specify that $\frac{1}{i\mathrm{th}\textit{PF}}$ and (i+1)thPF should be greater than 1,000. The results show 231 cases, only accounting for \$366.12. It suggests that this suspicious auction, which is overlooked by previous works, has been conducted on a small scale, and its collusive feature for the addresses involved should be taken into research.
@@
7.5. User behavior
Figure 9. The total number of ETH transfer events at different times before and after an NFT sale event.
In Unprofitable Trading, sellers collude with buyers to create unprofitable TXNs. For example, just 4 seconds before the sale event 0x1a4521, 0x4EDE98 completed a transfer (0.96 ETH) fully sufficient to support 0xE1f14d in buying Animeta #4302 (0.435 ETH) in a normal TXN 0x536fa8. To understand to what extent ETH transfer helps with each sale event, we define $M_{r}$ as the total amount received by sellers, $M_{t}$ as the total amount returned by sellers, and $M_{s}$ as the related NFT trading price. We then adopt Pearson correlation coefficient (Cohen et al., 2009) to measure the strength of the linear relationship between ($M_{r}$, $M_{s}$) and ($M_{t}$, $M_{s}$). The results are 0.7999 for ($M_{r}$, $M_{s}$) and 0.9840 for ($M_{t}$, $M_{s}$), indicating that the amount of each ETH transfer is similar to the NFT trading price. Furthermore, we count the number of ETH transfer events detected under a finer granularity (i.e., 60 seconds). Figure 9 shows that the closer the time point before/after a sale event is, the more ETH transfer behaviors happen. Also, wash traders make the most frequent ETH transfer between two and four minutes before/after a sale event, after which the number gradually decreases. The respective high price and time correlations prove that these ETH transfers are directly for the NFT TXNs.
@@ 7.6. NFT ecosystem
Table 5. Statistical results of decreased wash trading behavior
To better understand the impact of wash trading pairs/groups on NFT wash trading, we identify all 10,594 TXNs related to these address pairs and exclude their trading volume. For instance, if there is a flagged TXN in Hidden Trading, we do not calculate the trading volume of all consecutive private trading. As shown in Table 5, the results indicate a decrease of around 30% in wash trading amount for round-trip and Hidden Trading, while Unprofitable Trading experiences a significant decline, indicating the prevalence of wash traders in Unprofitable Trading. Interestingly, the addresses involved in wash trading pairs/groups accounted for only 0.082% of the total, meaning that a small group of users could cause considerable wash trading, and sizable wash trading behaviors could be eliminated by identifying and flagging them. It suggests that to avoid wash trading better and alert the users, NFT marketplaces do not need to keep all the trading addresses under governance but mark a small proportion of malicious accounts in event sequences.
@@
8. DISCUSSION
8.1. Threats to validity
Internal validity: 1) Adjusting the support for FP-Growth may affect the experimental results, e.g., the number of wash trading pairs/groups would increase if it is set loosely. Moreover, the support should be reconsidered for different datasets. To ensure the reliability of our results, we set a strict value, i.e., address pairs will be considered suspicious only when they participate in wash trading at least twelve times. 2) Our method may ignore unknown wash trading. We have not yet investigated more complex trading networks, e.g., incorporating internal TXNs to understand the intention of NFT traders from a more underlying perspective and thus reveal more NFT wash trading patterns. However, we are still the most comprehensive study on NFT wash trading.
External validity: The current tools (e.g., OpenSea API) for collecting NFT events have shortcomings, such as mislabeling transfer/sale/minted events. Nevertheless, we address the related issues and make public a guaranteed dataset for researchers.
8.2. Implications for NFT Marketplace